How many years will the following take?
$165 to grow to $46,967 if invested at 17.74 percent, compounded annually.
Answer #1

cash inflow*Present value of discounting factor(rate%,time period)

46,967=165*(1.1774)^n

(46,967/165)=(1.1774)^n

Taking log on both sides;

log (46,967/165)=n*log (1.1774)

n=log (46,967/165)/log (1.1774)

=34.60 years(Approx).